OpenStudy (anonymous):

write an equation of the line with the given slope and y-intercept. slope:3/4 y-intercepts:-6

jimthompson5910 (jim_thompson5910):

\[\Large y = mx+b\] \[\Large y = \frac{3}{4}x-6\]

OpenStudy (anonymous):

\[y = mx+b\]where m is the slope and b is the y-intercept.

OpenStudy (anonymous):

y= mx+b so, y=3/4x-6